# A Scalable Optimization Approach for Equitable Facility Location: Methodology and Transportation Applications

### v3

@@ 2.2.3 The Kolm-Pollak EDE
The Kolm-Pollak Equally-Distributed Equivalent (EDE) was introduced by Maguire and Sheriff (2020) based on Kolm–Pollak preferences Kolm (1976) and is uniquely suited for evaluating distributions of both amenities and burdens in transportation systems. Unlike previous measures, it satisfies all six properties required for equity assessment. However, like other nonlinear equity measures, its direct incorporation into optimization models has thus far been limited by computational complexity.
@@
$\mathcal{K}(\mathbf{z}):=-\frac{1}{\kappa}\ln\left[\frac{1}{N}\sum_{i=1}^{N}e^{-\kappa z_{i}}\right]$ (2)
The parameter $\kappa=\alpha\epsilon$ incorporates society’s aversion to inequality through $\epsilon$, where the sign of $\epsilon$ depends on whether larger or smaller values of the quantity are preferred. For desirable quantities like income or accessibility, where more is better, $\epsilon>0$ and the Kolm-Pollak EDE will be less than or equal to the mean. Conversely, for undesirable quantities like travel distance or exposure to pollution, where less is better, $\epsilon<0$ and the EDE will be greater than or equal to the mean. In either case, the magnitude of $|\epsilon|$ (typically between 0.5 and 2) represents the strength of aversion to inequality. Here $\alpha$ is a normalization factor, translating $\epsilon$ into $\kappa$:
$\alpha:=\frac{\sum_{i=1}^{N}z_{i}}{\sum_{i=1}^{N}z_{i}^{2}}$ (3)
Figure 1 illustrates this concept for a distribution of travel distances to facilities or resources. The histogram shows how distances are distributed across the population, with the mean distance indicated by a vertical line. Since travel distance is undesirable, we use $\epsilon<0$, resulting in an EDE that exceeds the mean. The gap between the EDE and mean represents the inequality penalty - larger values of $|\epsilon|$ increase this penalty, reflecting greater concern for residents experiencing particularly long travel times. This provides transportation planners with a single metric that captures both average system performance and distributional fairness.
A key advantage of the Kolm-Pollak EDE is how it bridges between conventional facility location objectives. As $\epsilon$ approaches 0, the EDE approaches the mean distance ($p$-median objective), while as $\epsilon$ approaches negative infinity, it approaches the maximum distance ($p$-center objective). This provides a principled way to balance average accessibility against the experience of the worst-off residents.
@@
Table 1 demonstrates these properties with four example distance distributions sharing the same mean but varying inequality. As disparity increases, the EDE grows larger than the mean, with the gap widening for larger values of $|\epsilon|$. This captures how inequitable access reduces overall system performance from an environmental justice perspective.
@@ 3.1 Base Model Formulation
Transportation planners and geographers need to evaluate how equitably a network serves its population, but existing metrics fail to capture both average performance and fairness in a single measure. The Kolm-Pollak EDE addresses this by quantifying accessibility while penalizing disparities in service. To apply this metric to facility networks, we measure the distance each resident must travel to reach their nearest facility, weighted by population density since residential data is typically aggregated at the census block level (the USA’s smallest census reporting area). Let $p_{r}$ represent the population of residential area $r\in R$, and let $z_{r}$ represent the distance that residents of $r$ must travel to reach a facility. The Kolm-Pollak EDE for this distribution is,
@@
where $T:=\sum_{r\in R}p_{r},$ the total population over all residential areas, and $\kappa=\alpha\epsilon$, where
$\alpha~{}=~{}\frac{\sum_{r\in R}p_{r}z_{r}}{\sum_{r\in R}p_{r}(z_{r})^{2}}.$ (5)
We model the distance traveled by residents of $r\in R$ as $z_{r}:=\sum_{s\in S}y_{r,s}d_{r,s}$ (where $y_{r,s}$ is as defined in Section 2), so we have
$\mathcal{K}(\mathbf{y})=-\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}p_{r}e^{-\kappa\sum_{s\in S}y_{r,s}d_{r,s}}\right].$ (6)
We can simplify this expression for use in an optimization model, which is a necessary step given the size of the instances we wish to solve.
The first thing to note is that $\alpha$, and therefore $\kappa=\alpha\epsilon$, depends on our optimization variables. However, we can greatly simplify the optimization by choosing to treat $\alpha$ as a constant. Fortunately, we can do this and still be minimizing an actual Kolm-Pollak EDE. Recall that $\alpha$ scales $\epsilon$, the chosen inequality aversion parameter, for use with the precise distribution we wish to evaluate. This means that the “true” value of $\alpha$ corresponds to the distribution given by an optimal solution to the optimization model, which we do not have a priori. However, we can estimate this distribution, and therefore $\alpha$, using the problem data. Crucially, we can quantify exactly how using an estimate for $\alpha$ impacts our solution. If $\alpha$ were expressed as a function of the decision variables in the objective function, then the aversion to inequality used to obtain the optimal solution would exactly match the $\epsilon$ selected by the user. The consequence of fixing $\alpha$ at the outset is that the user does not have complete control over the chosen level of aversion to inequality. For example, the user may have intended to choose an inequality aversion level of $\epsilon=-1$, however, because this is only scaled approximately, in reality, the solution obtained by the model optimizes a different level of inequality aversion. Once we have obtained this optimal solution, we can calculate the true value of $\epsilon$ and compare it to the level of inequality aversion that was intended. Given that the choice of inequality aversion is somewhat subjective, it may be that we are happy with the approximation. If not, we can resolve the model with an updated (more accurate) $\alpha$-approximation. See Section 3.3 for a detailed discussion.
@@
First recall that in our setting, $\epsilon<0$, and so we can remove the positive constant $-\frac{1}{\kappa}$ for optimizing $\mathcal{K}$. Then observe that because the natural logarithm function is monotonically increasing, we can equivalently minimize the argument of the function. Finally, dropping the positive constant $\frac{1}{T}$ gives us $\breve{\mathcal{K}}$. We formally define the Kolm-Pollak facility location model (KP) as follows, using the same notation as defined for (p-Med) above:
@@ 3.2.3 Split Demand Assignment
In some applications, it makes sense to allow the total demand of customer $r\in R$ to be served by multiple open facilities. Allowing for split demands is a natural consideration when service facilities have capacities, and it might not be possible for each customer to be assigned to the closest open facility. We allow for split demands by relaxing the integrality of the $y_{r,s}$ variables; i.e., by replacing equation (1f) with $y_{r,s}\in[0,1]$.
@@
Interestingly, when the $y_{r,s}$ variables are relaxed to the continuous interval $[0,1]$, Proposition 1 no longer holds, but (KPL) may actually be preferred over (KP) as the more accurate model. Recall that for a given $r\in R$, $\sum_{s\in S}y_{r,s}=1$ (constraint (1d)).
First consider the original, nonlinear Kolm-Pollak expression,
$-\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}p_{r}e^{-\kappa\sum_{s\in S}y_{r,s}d_{r,s}}\right].$
@@
For a fixed $r\in R$, the expression $\sum_{s\in S}y_{r,s}d_{r,s}$ represents a weighted average of the distances from $r$ to each of the facilities assigned to $r$. The distribution of distances measured by this Kolm-Pollak expression includes this weighted average distance for each of the $p_{r}$ people residing in census block $r$. Note that poor distances may be obscured in the weighted average of distances that is taken before the Kolm-Pollak expression is applied. This is not the case in the expression minimized by (KPL).
@@
We obtain the expression that is minimized by (KPL) by substituting the linear proxy into the Kolm-Pollak EDE:
$-\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}\sum_{s\in S}p_{r}y_{r,s}e^{-\kappa d_{r,s}}\right].$
@@
In this version of the Kolm-Pollak expression, the weighted average defined by $y_{r,s}$ for a given $r$ is applied to the collection of Kolm-Pollak terms of the form $p_{r}e^{-\kappa d_{r,s}}$, rather than to the distances directly. In other words, (KPL) models the scenario as if $p_{r}y_{r,s}$ people travel to facility $s$ at a distance of $d_{r,s}$. Because $p_{r}y_{r,s}$ may be fractional, the expression minimized by (KPL) is not technically a Kolm-Pollak score. However, this expression is a more accurate model of the Kolm-Pollak EDE in the case of split demands than the expression minimized by (KP), because distances are appropriately penalized.
@@
(KPL) with the $y_{r,s}$ variables relaxed has the same structure as (KPL) except with many binary variables replaced by continuous variables, so we would expect its computational performance to be at least as good as the computational performance of (KPL).
3.3 Parameter Calibration and Scaling
To minimize the Kolm-Pollak EDE via (KPL), we must approximate $\alpha$, the factor used to scale the inequality aversion parameter. As a consequence, we do not have complete control over the aversion to inequality represented by the optimal Kolm-Pollak score. To computationally explore the impact of approximating $\alpha$ (on both the aversion to inequality represented by the optimal distribution and on the optimal Kolm-Pollak score itself), we used five election polling scenarios. Data for the polling scenarios is summarized in Table 2. $|R|$ indicates the number of census blocks (residential areas), $|S|$ indicates the number of potential polling sites, and $k$ indicates the number of polling sites selected (which matches the number of polling sites used in the given election year). The potential polling sites have varying capacities so we use the (KPL) model with constraint (10). We set the desired aversion to inequality, denoted $\epsilon_{0}$, equal to $-2$ for all scenarios.
Scenario City Election $|R|$ $|S|$ $k$ Number of Variables
1 Salem 2016 3102 92 8 285,476
2 Richmond 2016 6208 211 65 1,310,099
3 Atlanta 2016 8364 387 126 3,237,255
4 Cincinnati 2016 7670 517 256 3,965,907
5 Baltimore 2020 9238 901 201 8,324,339
As a pre-processing step, we estimated the optimal distribution of distances by assigning each census block to the closest existing polling location. We used this distribution of distances to calculate an approximate value of $\alpha$, which we denote $\alpha_{a}^{in}$. Let (KPL)a represent the model (KPL) with $\kappa=\kappa_{a}:=\alpha_{a}^{in}\epsilon_{0}$. Let $\alpha_{a}^{out}$ denote the value of $\alpha$ associated with the optimal solution to (KPL)a. We can calculate the aversion to inequality represented by the optimal solution to (KPL)a as $\epsilon_{a}:=\kappa_{a}/\alpha_{a}^{out}$. Comparing $\epsilon_{0}$ and $\epsilon_{a}$ is a useful way to check the accuracy of $\alpha_{a}^{in}$. The closer $\alpha_{a}^{in}$ approximates $\alpha$, the closer $\epsilon_{a}$ will be to $\epsilon_{0}$. In Table 3, we see that $\epsilon_{a}$ ranges between $-1.56$ and $-0.55$. The desired level of aversion ($-2$) is not maintained by any of the optimal solutions.
Scenario $\epsilon_{0}$ $\epsilon_{a}$ $\epsilon_{\mathcal{K}}$ $\alpha_{a}^{in}$ $\alpha_{a}^{out}=\alpha_{\mathcal{K}}^{in}$ $\alpha_{\mathcal{K}}^{out}$
1 -2 -1.56 -2.00 0.000192 0.000246 0.000246
2 -2 -0.90 -2.00 0.000212 0.000474 0.000474
3 -2 -0.74 -1.98 0.000179 0.000485 0.000488
4 -2 -0.55 -2.01 0.000171 0.000616 0.000614
5 -2 -1.42 -2.01 0.001277 0.001797 0.001789
Table 3: Exact and approximate values of the aversion to inequality associated with the approximate solution where $\epsilon_{0}$ is the “desired” aversion to inequality.
It is reasonable to expect that the distribution of distances associated with an optimal solution to (KPL)a will produce a more accurate $\alpha$-estimate. Thus, we solve (KPL) again, this time with $\kappa=\kappa_{\mathcal{K}}:=\epsilon_{0}\alpha_{\mathcal{K}}^{in}$, where $\alpha_{\mathcal{K}}^{in}:=\alpha_{a}^{out}$. We denote this version of (KPL) as (KPL)K. We compute the aversion to inequality associated with the optimal solution to (KPL)K as $\epsilon_{\mathcal{K}}:=\kappa_{\mathcal{K}}/\alpha_{\mathcal{K}}^{out}$, where $\alpha_{\mathcal{K}}^{out}$ is calculated using the optimal distribution of distances. In Table 3, we see that $\epsilon_{\mathcal{K}}$ is very close to -2 for every scenario: as expected, $\alpha_{\mathcal{K}}^{out}$ provides a very good approximation to $\alpha$.
Due to the high quality of its $\alpha$-approximation, (KPL)K generates high-quality solutions in the sense that the desired level of aversion to inequality is very nearly preserved by its optimal solutions. Thus, we used optimal solutions to (KPL)K as a benchmark to test the quality of optimal solutions to (KPL)a. In Table 4, we compared the Kolm-Pollak EDE values associated with optimal solutions to (KPL)a and (KPL)K. The third column, labeled $|\Delta|$, reports the absolute difference between the optimal Kolm-Pollak EDEs (in meters). The values in the last column, labeled “gap”, are found by dividing $\Delta$ by the Kolm-Pollak EDE associated with (KPL)K. For each of these five scenarios, (KPL)a produced near-optimal solutions.
KP EDE
Scenario (KPL)a (KPL)k $|\Delta|$ gap
1 6541.81 6541.81 0.00 0.0000
2 2256.18 2259.49 3.31 0.0015
3 2198.29 2191.37 6.92 0.0032
4 5585.93 5571.38 14.54 0.0026
5 571.21 566.15 5.06 0.0089
Table 4: KP EDEs obtained via (KPL)a are close to the best-know KP EDEs, which were obtained via (KPL)K. Distances are measured in meters.
We chose to use polling scenarios for this analysis (as opposed to the food desert study described in Section 4) because four of the polling scenarios optimally locate many more than 10 locations, and the polling models allow all existing sites to relocate. This means that the $\alpha$-approximations will be less accurate than in the food desert application in which we assume existing supermarkets remain open and the maximum value of $k$ is 10. In fact, the Kolm-Pollak score improved by a mile or more in every polling instance, so the distribution of distances before and after optimization was quite different. Although small, this study provides some assurance that optimal solutions are not highly sensitive to minor variations in $\alpha$, or equivalently, to varying levels of aversion to inequality.
Despite this, some modelers may wish to have more control over the aversion to inequality represented by the optimal solution, and our experiments point to the following strategy. Given a desired aversion to inequality, $\epsilon_{0}$, solve (KPL) using a value of $\alpha$ that is estimated using problem data, $\alpha_{in}$. Next, calculate the aversion to inequality represented by the optimal distribution of distances: $\epsilon_{out}=\frac{\alpha_{in}\epsilon_{0}}{\alpha_{out}}$, where $\alpha_{out}$ is calculated using the optimal distribution of distances. If the user is not happy with the aversion to inequality that the solution represents, they may re-solve the model with the more accurate $\alpha$-approximation, $\alpha_{out}$. According to our experiments, the next optimal solution is likely to correspond to an aversion to inequality that is very close to $\epsilon_{0}$.

### v4

@@ 2.2.3 The Kolm-Pollak EDE
The Kolm-Pollak Equally-Distributed Equivalent (EDE) was introduced by Maguire and Sheriff (2020) based on Kolm–Pollak preferences (Kolm, 1976), and is uniquely suited for evaluating distributions of both amenities and burdens in transportation systems. Unlike previous measures, it satisfies all six properties required for equity assessment.
@@
$\mathcal{K}(\mathbf{z}):=-\frac{1}{\kappa}\ln\left[\frac{1}{N}\sum_{i=1}^{N}e^{-\kappa z_{i}}\right].$ (2)
The parameter $\kappa:=\alpha\epsilon$ incorporates society’s aversion to inequality through $\epsilon$, where the sign of $\epsilon$ depends on whether larger or smaller values of the quantity are preferred. For desirable quantities like income or accessibility, where more is better, $\epsilon>0$ and the Kolm-Pollak EDE will be less than or equal to the mean. Conversely, for undesirable quantities like travel distance or exposure to pollution, where less is better, $\epsilon<0$ and the EDE will be greater than or equal to the mean. In either case, the magnitude $|\epsilon|$ (typically between 0.5 and 2), represents the strength of aversion to inequality. Here $\alpha$ is a normalization factor, translating $\epsilon$ into $\kappa$:
$\alpha:=\frac{\sum_{i=1}^{N}z_{i}}{\sum_{i=1}^{N}z_{i}^{2}}.$ (3)
Figure 1 illustrates this concept for a distribution of travel distances to facilities or resources. The histogram shows how distances are distributed across the population, with the mean distance indicated by a vertical line. Since travel distance is undesirable, we use $\epsilon<0$ resulting in an EDE that exceeds the mean. The gap between the EDE and mean represents the inequality penalty – larger values of $|\epsilon|$ increase this penalty, reflecting greater concern for residents experiencing particularly long travel times. The EDE provides transportation planners with a single metric that captures both average system performance and distributional fairness.
@@
In addition to satisfying the required properties, a key advantage of the Kolm-Pollak EDE is how it bridges between conventional facility location objectives. As $\epsilon$ approaches 0, the EDE approaches the mean distance ($p$-median objective), while as $\epsilon$ approaches negative infinity, it approaches the maximum distance ($p$-center objective). This provides a principled way to balance average accessibility against the experience of the worst-off residents. Moreover $\epsilon$, the so-called “inequality aversion” parameter, is consistent with other EDE functions from the literature, for example the Atkinson EDE, and is well-understood by the environmental justice community.
@@
Table 2 demonstrates properties of the Kolm-Pollak EDE using four example distance distributions sharing the same mean but having varying inequality. As disparity increases, the EDE grows larger than the mean, with the gap widening for larger values of $|\epsilon|$. This captures how inequitable access reduces overall system performance from an environmental justice perspective.
@@ 3.1 Base Model Formulation
Transportation planners and geographers need to evaluate how equitably a network serves its population. The Kolm-Pollak EDE addresses this by quantifying accessibility while penalizing disparities in service. To apply this metric to facility networks, we measure the distance each resident must travel to reach their nearest facility, weighted by population density since residential data is typically aggregated at the census block level (the USA’s smallest census reporting area). Let $p_{r}$ represent the population of residential area $r\in R$, and let $z_{r}$ represent the distance that residents of $r$ must travel to reach a facility. The Kolm-Pollak EDE for this distribution is,
@@
where $\mathbf{z}\in\mathbb{R}^{|R|}$ is the vector of travel distances, $T:=\sum_{r\in R}p_{r}$ is the total population over all residential areas, $\kappa:=\alpha\epsilon$, and
$\alpha~:=~\frac{\sum_{r\in R}p_{r}z_{r}}{\sum_{r\in R}p_{r}(z_{r})^{2}}.$ (5)
We model the distance traveled by residents of $r\in R$ as $z_{r}:=\sum_{s\in S}y_{r,s}d_{r,s}$ (where $y_{r,s}$ and $d_{r,s}$ are as defined in Section 2), so we have
$\mathcal{K}(\mathbf{y})=-\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}p_{r}e^{-\kappa\sum_{s\in S}y_{r,s}d_{r,s}}\right],$ (6)
with $\mathbf{y}\in\mathbb{R}^{|R|\times|S|}$. We can simplify this expression for use in an optimization model, which is a necessary step given the size of the instances we wish to solve.
The first thing to note is that $\alpha$, and therefore $\kappa=\alpha\epsilon$, depends on our optimization variables. However, we can greatly simplify the optimization by choosing to treat $\alpha$ as a constant. Fortunately, we can do this and still be minimizing an actual Kolm-Pollak EDE.
Recall that $\alpha$ scales $\epsilon$, the chosen inequality aversion parameter, for use with the precise distribution we wish to evaluate. This means that the “true” value of $\alpha$ corresponds to the distribution given by an optimal solution to the optimization model, which we do not have a priori. However, we can estimate this distribution, and therefore $\alpha$, using the problem data.
Crucially, we can quantify exactly how using an estimate for $\alpha$ impacts our solution. If $\alpha$ were expressed as a function of the decision variables in the objective function, then the aversion to inequality used to obtain the optimal solution would exactly match the $\epsilon$ selected by the user. The consequence of fixing $\alpha$ at the outset is that the user does not have complete control over the chosen level of aversion to inequality. For example, the user may have intended to choose an inequality aversion level of $\epsilon=-1$, however, because this is only scaled approximately, in reality, the solution obtained by the model optimizes a different level of inequality aversion. Once we have obtained this optimal solution, we can calculate the true value of $\epsilon$ and compare it to the level of inequality aversion that was intended. Given that the choice of inequality aversion is somewhat subjective, it may be that we are happy with the approximation. If not, we can re-solve the model with a more accurate $\alpha$-approximation. Extensive computational tests yielded a near-perfect level of inequality aversion on the second run of the model in every instance. See Section 4 for a detailed discussion.
@@
First recall that in our setting, $\epsilon<0$, and so we can remove the positive constant $\frac{1}{\kappa}$ for optimizing $\mathcal{K}$. Then observe that because the natural logarithm function is monotonically increasing, we can equivalently minimize the argument of the function. Finally, dropping the positive constant $\frac{1}{T}$ gives us $\breve{\mathcal{K}}$. We formally define the Kolm-Pollak facility location model (KP) as follows, using the same notation as defined for (p-Med) above:
@@ 3.2.3 Split Demand Assignment
In some applications, it makes sense to allow the total demand of customer $r\in R$ to be served by multiple open facilities. Allowing for split demands is a natural consideration when service facilities have capacities, and it might not be possible for each customer to be assigned to the closest open facility. We allow for split demands by relaxing the integrality of the $y_{r,s}$ variables; i.e., by replacing constraints (1f) with $y_{r,s}\in[0,1]$.
@@
Interestingly, when the $y_{r,s}$ variables are relaxed to the continuous interval $[0,1]$, Proposition 1 no longer holds, but (KPL) may actually be preferred over (KP) as the more accurate model. Recall that for a given $r\in R$, constraints (1d) ensure $\sum_{s\in S}y_{r,s}=1$, and first consider the original, nonlinear Kolm-Pollak expression,
$\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}p_{r}e^{-\kappa\sum_{s\in S}y_{r,s}d_{r,s}}\right].$ (6)
@@
For a fixed $r\in R$, the expression $\sum_{s\in S}y_{r,s}d_{r,s}$ represents a weighted average of the distances from $r$ to each of the facilities assigned to $r$. The distribution of distances measured by expression (6) assumes this weighted average distance for each of the $p_{r}$ people residing in census block $r$, and as such, inequality between residents in the same census block is not accounted for in the Kolm Pollak score. Poor distances are obscured because the weighted average is taken before the Kolm-Pollak expression is applied. However, this is not the case in the expression minimized by (KPL).
@@
Substituting the linear proxy into the Kolm-Pollak EDE, we obtain the expression that is minimized by (KPL):
$\frac{1}{\kappa}\ln\left[\frac{1}{T}\sum_{r\in R}\sum_{s\in S}p_{r}y_{r,s}e^{-\kappa d_{r,s}}\right].$ (11)
@@
In this version of the Kolm-Pollak expression, the weighted average defined by $y_{r,s}$ for a given $r$ is applied to the collection of Kolm-Pollak terms of the form $p_{r}e^{-\kappa d_{r,s}}$, as opposed to the distances directly. In other words, (KPL) models the scenario as if $p_{r}y_{r,s}$ people travel to facility $s$ at a distance of $d_{r,s}$, and therefore does account for any inequality between residents in the same census block. Note that because $p_{r}y_{r,s}$ may be fractional, the expression minimized by (KPL) is not technically a Kolm-Pollak score. However, the expression is a more accurate model of the Kolm-Pollak EDE any time we wish to take into account the inequality within a given residential area because distances are appropriately penalized. See Example 2 for a concrete illustration of these concepts.
Example 2.
Consider a single residential area with a total population of 100. Suppose half of the population are assigned to a facility located at a distance of 1 unit away and half are assigned to a facility located 100 units away as illustrated in Figure 2.
Figure 2: Illustration of the distribution in Example 2.
For this distribution, we compute
$\alpha=\frac{(50\times 1)+(50\times 100)}{(50\times 1^{2})+(50\times 100^{2})}\approx 0.0101,$
and setting $\epsilon=-1$, obtain $\kappa=-0.0101.$
Using these values, we substitute into expression (6), and calculate the Kolm-Pollack score of the distribution to be
$\frac{1}{\kappa}\ln\left[\frac{1}{100}\times 100\times e^{-\kappa\Big(\big(\frac{1}{2}\times 1\big)+\big(\frac{1}{2}\times 100\big)\Big)}\right]=50.5.$
Observe that the value obtained is precisely the mean of the distribution. Moreover, this occurs because the model assumes that each person residing at location A travels the same distance to reach a facility, at least on average. As there is only one residential area in this simple example much of the expression cancels, and the Kolm Pollak score reduces to the mean. This is expected because with only one residential area under consideration, there can be no inequality between different residential areas, and inequality within the same residential area is not accounted for when using expression (6).
In contrast, substituting into equation (11), we obtain a value strictly greater than the mean, accounting for the inequality present within our single residential area:
$\frac{1}{\kappa}\ln\left[\frac{1}{100}\times 100\times\Bigg(\left(\frac{1}{2}\times e^{-\kappa\times 1}\right)+\left(\frac{1}{2}\times e^{-\kappa\times 100}\right)\Bigg)\right]\approx 62.39.$
Using expression (11), we are able to model the inequality in the proposed scenario by making the implicit assumption that 50 people travel a distance of 1 to reach a facility, and 50 people travel a distance of 100 to reach a facility.
@@
Note that (KPL) with the $y_{r,s}$ variables relaxed has the same structure as (KPL) except with many binary variables replaced by continuous variables, so we would expect its computational performance to be at least as good as the computational performance of (KPL).